Help!
Find the length of the diagonal of a cube with side length 10 cm.

Answer by RAY100(1637)

diagonal ^2 = l^2 +w^2 +z^2
.
for a cube,, d^2 = 3 s^2.
.
for s =10, d^2 = 3 (10)^2 =300
.
dia = sqrt 300 = 17.32
.